Please help me with this problem:
Working together, Ann and Vicky can finish the report in 2 2/9 hours. If Ann can type a report in 5 hours, how many hours can Vicky type the same report alone?

Answer by mananth(16946)

Ann & Vicky can finish report in 2 2/9 hours = 20/9 hours
so they can do 9/20 of the job in 1 hour

Ann can do the same job in 5 hours
so she does 1/5 of the job in 1 hour.
1/5 - 9/20 = job done by Vicky in 1 hour
LCD = 20
(20-9)/20
11/20
so he takes 20/11 hours to do the job alone
1 9/11 hours